OpenStudy (anonymous):

What is the prime factorization of 560?

OpenStudy (tkhunny):

You don't need to be in a hurry to find them. You also needn't find them all at once. Just find ANY factor and see where it leads. 560 is even. There is a factor of 2 560 = 2*280 280 is even, there is a factor of 2 \(560 = 2*2*140 = 2^{2}\cdot 140\) Continue on this way until you have only prime factors remaining.

OpenStudy (tkhunny):

It doesn't always have to be 2. Whatever you see is good. Maybe next would be... \(560 = 2^{2}\cdot 140 = 2^{2}\cdot 10\cdot 14\) Not this time, but sometimes, you need to know that 51 = 3*17 :-) People mistake 51 for prime. Nope. You have to have the Prime Numbers in your head.